Report of Activities on the RVIB N.B. Palmer Cruise 02-04
On 22 August, we were at the seaward end of survey line 7 finishing up work at station 45 that was started on the 21st. A second Tucker trawl was completed around 0130. This tow, while completed successfully, was not without its problems. A slowdown of the ship when it steamed through some heavier pack ice caused the cable towing angle to drop rapidly and, with more wire out than water depth, the net grazed the sea floor. When it returned to the deck, some benthic animals in addition to planktonic ones were in the catch. A 1-m MOCNESS tow to a 1000 m followed and it too had difficulties. While the unit was still deep and coming back to the surface the towing wire snagged on a large chunk of pack ice flowing into the wake region. The sudden release of the wire when it broke free of the ice caused the wire to loop over a stanchion welded onto the railing on the port side of the stern. It took some time and some clever maneuvering of the ship before the wire was unhooked and the towing resumed. With the repairs to the ROV completed, an ROV under-ice survey was the final activity at station 45. BIOMAPER-II was deployed at the start of the steam to station 46, which was located on the edge of the continental shelf.
The twenty-two mile steam to
Station 46 took approximately 9 hours. As we approached the station, we were
amazed to see the ice buoy station that had been installed on the ice flow
where the Gould had established their
first process station between stations 75 and 76. The seabird observers spotted
the ice station while doing their surveying.
Along with the ice station, which has satellite telemetry to beam the
data back to a shore station in the
The work at station 46 consisted of 2 CTD casts, and three Tucker trawls to collect live animals. There was an intention to do an under-ice SCUBA dive, but this was scratched because of the cold temperatures; below about -18ºC regulators can freeze up before the divers can enter the water. As noted below, the frigid air quickly froze some of the sensors on the CTD and required special tactics to get them into water still functioning. The second Tucker trawl succumbed to a too quick descent once in the water and the cod-end of the net came up wrapped over the top net bar and there was no appreciable catch. The work was completed in about 5 hours and the steam to station 64 commenced in the evening. BIOMAPER-II was delayed going into the water for the transit to station 64 until it was determined that the pack ice conditions were suitable for towing.
The weather was basically a repeat of yesterday's and we remained in the deep freeze. Temperatures varied from -24ºC in the early morning to -27ºC in the late evening. The barometer began a slow climb during the day from 985 to 988 mb. Winds were mostly in the kt range out of the southwest. Consequently, the working conditions on deck were OK in spite of the cold. The pack ice was mostly 9/10 or 10/10. Skies were crystal clear and nearly cloudless all day. Again the combination of a setting sun and a rising full moon provided a spectacular close to the day.
CTD Group report (Eileen Hofmann,
Bob Beardsley, Baris Salihoglu,
Chris MacKay, Francisco (
On 22 August, we completed one CTD cast at station 46, which is at the outer end of survey transect 7. The station consisted of a short cast to 100 m for the FRRF profile. Following this cast, the CTD was lowered to within a few meters of the bottom, which was at about 438 m.
This cast showed what the incredibly cold air temperatures (-21ºC to -23ºC and -40ºC to -41ºC with the wind chill) can do to the CTD sensors. In the short time it took to get the CTD out the Baltic Room door and lowered into the water, the water in the tubing connecting the various CTD sensors froze. As a result the conductivity and oxygen readings obtained as the CTD was lowered were in error. Therefore, after the short FRRF cast, the CTD was brought back into the Baltic Room, the water in the tubing was thawed, and saltwater was pumped through the sensors and into the tubing. The CTD was returned to the water and the second cast was done to the bottom. This cast returned good data, in spite of an abbreviated soak time at the surface to equilibrate the conductivity and oxygen sensors.
The temperature profile from this station showed a well-mixed surface layer, at the freezing point, in the upper 100 m. Below this, temperature increased, reaching a maximum of 1.73ºC at about 250 m. Between 200 m and 350 m the temperature remained at 1.60 to 1.73ºC. Salinity at these depths was relatively constant at 34.66. This is consistent with the presence of Circumpolar Deep Water, which appears to be moving onto the continental shelf at this location.
The CTD cast at station 46 afforded those in the Baltic Room the opportunity to see a beautiful, extended Antarctic sunset and a lovely moon rise. The colors of the sky and reflected light from the sea ice were spectacular.
Sea Birds (Chris Ribic and Erik Chapman)
Surveys were conducted for 7 hours and 38 minutes on 22 August as the ship moved between stations 45 and 46. The recent cold, clear weather pattern continued today. Once again, vast floes covered 9 to 10/10ths of the ocean's surface. New gray and new white ice covered areas that were probably leads only days ago, prior to the recent cold weather. Snow Petrels were uncommon, but occasionally present in the survey. It appeared that some of the Snow Petrels were moving in the direction of the ice-edge when they turned to check out the broken ice behind the ship. Throughout most of the day, there were 2 to 3 Snow Petrels behind the ship, flying low over the broken ice and occasionally landing and pecking into the ice. As the ship stopped at station 46 in a recently frozen lead, 18 Crabeater Seals were observed off the port side of the ship.
A summary of the birds and marine mammals observed on 22 August (YD 234) during 7 hours, 38 minutes of survey time as the ship traveled between stations 45 and 46 is the following:
Species (common name)
Species (scientific name)
MOCNESS/ADCP/OPC Report (Phil Alatalo, Ryan Dorland, Peter Wiebe, Dicky Allison, Scott Gallager, Gareth Lawson)
The 1-m MOCNESS was deployed at station 45 (-67º 11.22′S; -74º 27.19′W) at 0200 on 22 Aug. It was a deep 1000 m slope-water tow. Unfortunately, the thick ice caused an unusual problem while fishing net 3. The tow wire caught on ice as the ship turned. When freed, the wire looped around a vertical stanchion located on the port side of the stern. Coordinated work by the bridge, aft-control winch operator, and the MT crew freed the wire without damage after about 25 minutes. During the interval, the MOCNESS plummeted down to 1000 m, remaining vertical for several minutes. The volume filtered data for this net remains questionable, since only one adult krill (and very little else) was present in the sample. The OPC functioned well until 200 m, when it lost communication with its deck box. The ice-related difficulties and communication problems between the OPC and the MOCNESS during tow 7 resulted in a loss of some OPC data. Retrieval was hampered by the extreme cold (-23.4ºC).
Taxonomic composition for this tow was unique to its offshore location: fish, jellyfish, and krill (Thysanoessa sp.) comprised most of the biomass between 1000 and 200 m. Above 200 m, krill and salps made up the largest portion of biomass. However, overall biomass collected in all nets was not very great given the large volumes filtered. Of interest were a large fish (Borastomias antarcticus) and 8-cm diameter jellyfish present in the deepest net. A myctophid fish, probably Gymnoscopelus braueri, was caught between 400 and 200 m and two unusual, dark brown pteropods were found below 400 m. Radiolarians, chaetognaths, siphonphores, and copepods were most numerous between 600 and 400 m. At the depth interval of 100-50 m, ostracods joined this assemblage.
Results from OPC data processed
from 1-m MOCNESS tow 6 taken at station 42 showed low counts and biovolume throughout the 384 m water column, with count
densities reaching 700 individuals/m3 around 300 m and decreasing to around 350
near the surface. In comparison, values in
BIOMAPER II group report (Gareth Lawson, Peter Wiebe, Scott Gallager, Phil Alatalo Dicky Allison, Alec Scott)
On August 22, we managed two BIOMAPER II towyos, the first of which saw us transiting across the continental shelf break from the deep station 45 (2966 m) to the more shallow station 46 (433 m). Despite strong noise affecting our up-looking transducers on those occasions where the ship was backing and ramming its way through particularly thick ice, we were able to make some interesting acoustic and video observations. For the whole length of the tow, a diffuse scattering layer was present between approximately 50 and 150 m, associated with the pycnocline as usual. On one occasion this layer was disrupted and replaced by a number of stronger single targets, only to reform a half mile later. With the VPR, we again made observations in this depth range of copepods, radiolarians, and diatoms. Near the start of the transect, we recorded a number of dense krill-like patches at about 250 m depth; unusually, these aggregations were very tall (60 m) but only 50 to 150 m in length. Just before we were forced to pull the tow-body up due to severe ice conditions, the bottom shoaled up sufficiently to reveal a diffuse scattering layer extending from the bottom to 100 m off-bottom.
Our second towyo took us back out across the continental shelf break between stations 46 and 64. The shallow scattering layer at the pycnocline was still present, though somewhat reduced in intensity from the previous tow. On occasion, thinner and denser sub-layers were evident within this overall layer, which the VPR indicated may have been composed of krill smaller than 1 cm (i.e., furcilia or other larval stages). Below 200 m, the VPR captured images of a diverse species assemblage, including copepods, gelatinous zooplankton, worms, and radiolarians. In addition, a series of fairly dense patches formed a nearly continuous acoustic layer at 350 m. Since we were in deep waters (800-3000 m) off the continental shelf where there was little risk of running the BIOMAPER II into the bottom, we decided to try to send the tow-body into this layer to capture images of its denizens with the VPR. The deepest point the BIOMAPER II reached was 390 m, where the VPR observed tomopterid worms and small copepods.
Current Position and Conditions
The offshore stations in the central sector of the survey grid have now been completed and the next effort will be to attempt to run in on survey line 8 as close to shore as the pack ice will permit the Palmer to penetrate in a reasonable length of time. We are now completing a CTD cast at station 63 and will shortly be headed for station 62. Our position at 2348 on 23 August is -67º 40.284′S; -74º 35.079′W. The air temperature is -23.1ºC and the barometric pressure is 994.8 mb. Winds are around 6 kts out of the south-southwest (214). High cirrus clouds are overhead, but the full moon and some stars are still visible.